OpenStudy (anonymous):

Cross-Multiply! 8 over 11 = 3 over 2y i got 19y=33 @mathmale

OpenStudy (mathmale):

Again i'd make a point of writing the equation of ratios as fractions, using ---- instead of /:\[\frac{ 8 }{ 11 }=\frac{ 3 }{ 2y }\] because it's easier to see what's happening when the equation is written in this manner. Cross multiply. Solve for y.

OpenStudy (anonymous):

19y=33 ..thats where im stuck

OpenStudy (mathmale):

You can divide both sides of the equation by 19, to isolate y, and then you'll have your answer. No need to go any further. Just leave your answer as "y = whatever."

OpenStudy (mathmale):

However, Taylor, you may want to go back and check your multiplication. What is 8(2y)?

OpenStudy (anonymous):

omg... 16y=33

OpenStudy (texaschic101):

correct....now divide both sides by 16

OpenStudy (mathmale):

Yes, and when you've done that, and have a fraction on the right side as the value of y, stop. You're done!

OpenStudy (anonymous):

2.06 is what i got. that is one of the answers on my wksht

OpenStudy (anonymous):

*isnt one of the answers

OpenStudy (texaschic101):

is 33/16 one of the answers

OpenStudy (mathmale):

that's because there's no need to divide. As I've suggested, leave your answer as a fraction, like this: y = a/b

OpenStudy (texaschic101):

or 2 1/16

OpenStudy (mathmale):

It's actually better that you DON'T divide, because you'd have to round off your answer (as you have, to 2.06) otherwise.

OpenStudy (mathmale):

In order from best answer to poorest: y=33/16, y=33 1/16, y= 2.06

OpenStudy (texaschic101):

or y = 2 1/16

OpenStudy (anonymous):

33/16 is an answer

OpenStudy (mathmale):

@texaschic101: Thanks for the correction: How embarrassing!

OpenStudy (texaschic101):

good job explaining mathmale

OpenStudy (mathmale):

Taylor: go for it.

OpenStudy (anonymous):

thanks for your help!

OpenStudy (texaschic101):

thats what we are here for :)